### v1

@@ 4.1.1 Properties of $\mathsf{PN}$
To begin, we recall one of the integral tools [Ren22] uses to study $\mathsf{PN}$ matchings. This is the notion of a (Reny-)feasible chain, which is defined as follows:
@@
A chain $\mu_{1}<\mu_{2}<\ldots<\mu_{N}$ is Reny-feasible if $\mu_{1}$ is stable, $\mu_{N}$ is Pareto-efficient, and for each $i$, $\mu_{i}$ does not violate the priority of any applicant who is Pareto-improvable in $\mu_{i-1}$.
While [Ren22] reasons about Reny-feasible sequences in quite some detail, he does not prove the following fundamental fact about Reny-feasible sequences: every matching in a Reny-feasible sequence is always priority-neutral. To prove this fact, we extend the techniques from [Ren22, Lemmas 3–6]. First, we recall another simple lemma from [Ren22]:
@@
We now prove our core result regarding $\mathsf{PN}$.
@@
Fix some $n\in\{1,\ldots,N\}$, and suppose that $\xi$ is an improver-only violator of $\mu_{n}$. Using induction on $n^{\prime}$, we will prove the following claim: for all $n^{\prime}\in\{1,\ldots,n\}$, we have that for all applicants $d$,111111 [Ren22, Lemmas 4 and 5] perform related inductive arguments, but get weaker conclusions that are only sufficient to conclude that the final matching $\mu_{N}$ in the sequence is priority-neutral.
@@
To begin, we will prove using 4.2 that $\mu_{n^{\prime}}\vee\xi$ is a matching. To see this, observe that by induction, we know that for all $d\in\operatorname*{vio}(\mu_{n^{\prime}})$, we have $\mu_{n^{\prime}}(d)\succeq_{d}\xi(d)$. Thus, $\mu_{n^{\prime}}$ does not violate any $d$’s priority at $\xi(d)$. On the other hand, by the definition of an improver-only violator of $\mu_{n}$, we know that for all $d\in\operatorname*{vio}(\xi)$, we have $\xi(d)\succeq_{d}\mu_{n}(d)\succeq_{d}\mu_{n^{\prime}}(d)$. Thus, $\xi$ does not violate any $d$’s priority at $\mu_{n^{\prime}}(d)$. Thus, 4.2 implies that $\mu_{n^{\prime}}\vee\xi$ is a matching.
Now, consider any applicant $d\in\operatorname*{vio}(\mu_{n^{\prime}+1})$. By the definition of a Reny-feasible sequence, $d$ is not Pareto-improvable in $\mu_{n}^{\prime}$. Since $\mu_{n^{\prime}+1}\geq\mu_{n^{\prime}}$, also by the definition of a Reny-feasible sequence, we thus have $\mu_{n^{\prime}+1}(d)=\mu_{n^{\prime}}(d)$. Thus, since $\mu_{n^{\prime}}\vee\xi\geq\mu_{n^{\prime}}$, we must have $[\mu_{n^{\prime}}\vee\xi](d)=\mu_{n^{\prime}}(d)$, and therefore $\mu_{n^{\prime}+1}(d)=\mu_{n^{\prime}}(d)\succeq_{d}\xi(d)$. This finishes the proof of the inductive step, and proves the claim for all $n^{\prime}\leq n$.
@@ 4.1.2 Properties of $\mathsf{L}$
Now, we turn to the legal lattice $\mathsf{L}$. To begin, recall that [EM20] call a set of matchings $\mathsf{L}$ legal if $\mathsf{L}=\{\mu\ \allowbreak|\ \allowbreak\text{$\forall\nu\in\mathsf{L}:$ \allowbreak no applicant $d$'s priority is violated by $\mu$ at institution $\nu(d)$}\}$. [EM20] prove that there exists a unique legal set $\mathsf{L}$, and that legal matchings also form a lattice which contains all stable matchings, and whose applicant-optimal element equals $\mathsf{EADAM}$. Moreover, [Ren22] shows that $\mathsf{PN}\subseteq\mathsf{L}$.
Unlike priority-neutral lattices, however, legal lattices always have the same structural properties as stable lattices. [FZ22] formalizes this by showing that every legal lattice can be represented as the stable lattice on some altered set of preferences and priorities; we make heavy use of this idea, and other techniques from [FZ22], below. In summary, the main known results are the following:
Theorem 4.4 ([EM20, Ren22, FZ22]).
Now, we introduce another type of chain in order to additionally reason about the legal lattice while bringing in the techniques of [TY14]. This chain modifies the Kesten-Tang-Yu chain (as defined in 2.3) by, informally speaking, progressing the chain with “as small of differences as possible.” To formalize this, we recall that in any lattice, an element $x$ covers $y$ if $x\geq y$, and for any $x\geq z\geq y$, we have either $x=z$ or $y=z$. Covering elements in stable lattice correspond to two matchings differing by rotations in the rotation poset. In other words, $\mu_{2}$ covers $\mu_{1}$ in a stable lattice if and only if there is a rotation $\alpha$ exposed in $\mu_{1}$ such that $\alpha\mu_{1}=\mu_{2}$ (by the notation of Section 2.1). See [GI89] (where covering relationships are discussed under the term “minimal differences”) or [CT19] for details.
@@
We say that a chain of matchings is a gradual Kesten-Tang-Yu chain if it can be constructed as follows: Set $P_{0}=P$ to be the full preference (and priority) profile, and let $\mu_{0}=\mathsf{IPDA}(P_{0})$. Now, for $i\geq 1$, as long as $\mu_{i-1}\neq\mathsf{APDA}(P_{i-1})$, set $P_{i}=P_{i-1}$, and set $\mu_{i}$ equal to any matching which covers $\mu_{i-1}$ in the stable matching lattice under preferences $P_{i-1}$. If $\mu_{i-1}=\mathsf{APDA}(P_{i-1})$, yet $\mu_{i-1}$ is not Pareto-optimal, then modify $P_{i}$ as in 2.3 (i.e., let $P_{i}$ denote the following operation on $P_{i-1}$: for each $d$ matched to an under-demanded institution in $\mu_{i-1}$, remove all institutions in their list strictly before $\mu_{i-1}(d)$). One can show $\mu_{i-1}$ is stable under preferences $P_{i}$ as well; we set $\mu_{i}$ to any matching stable under $P_{i}$ which covers $\mu_{i-1}$ in the stable lattice under $P_{i}$. The chain terminates with $\mu_{N}$ once $\mu_{N}$ is Pareto-optimal.
By [Ren22, Lemma 2], or by the techniques of [TY14], we know that if $\mu_{i}$ in the gradual Kesten-Tang-Yu chain is not Pareto optimal, then some applicants must be matched to under-demanded institutions.121212 In more detail, [Ren22, Lemma 2] shows that if an applicant $d$ is Pareto-improvable at $\mu_{i}=\mathsf{APDA}(P_{i})$, then $d$ cannot be matched to an under-demanded in $\mathsf{APDA}(P_{i})$. Thus, whenever $\mu_{i}=\mathsf{APDA}(P_{i})$ but $\mu_{i}$ is not Pareto-optimal, we know that $P_{i+1}\neq P_{i}$. Conversely, if the matching is Pareto-optimal, then it must equal $\mathsf{APDA}$ under the current preferences (and under the original preferences). Hence there always exists at least one gradual Kesten-Tang-Yu sequence.
Additionally, any gradual Kesten-Tang-Yu sequence is also Reny-feasible. To see why, observe that, by construction, a gradual Kesten-Tang-Yu sequence will only violated the priority of applicants matched to under-demanded institutions. Moreover, the first matching $\mu_{0}$ of a gradual Kesten-Tang-Yu chain equals $\mathsf{IPDA}$, and the last matching $\mu_{N}$ is Pareto-efficient, by definition. Hence, by 4.3, all elements of any gradual Kesten-Tang-Yu sequence are priority-neutral, and by Theorem 4.4, also legal.
@@
Fix some profile $P$ of preferences (and priorities). To begin, we recall some concepts from [FZ22]. If applicant $d$ is never matched to institution $h$ in any matching in $\mathsf{L}$, then [FZ22] call the pair $(d,h)$ an illegal pair. Let $\widetilde{P}$ denotes the altered preferences where, for all $d$, we modify $d$’s preference list by removing all $h$ such that $(d,h)$ is an illegal pair. [FZ22, Lemma 3.5] shows that the legal lattice on $P$ equals the stable lattice on $\widetilde{P}$ (and moreover that illegal pairs can always be removed from applicants preference lists without changing the set of legal matchings). Our proof proceeds by showing that $\widetilde{P}$ can be constructed from the execution of a gradual Kesten-Tang-Yu algorithm as in 4.5 (along with additional steps after the sequence) in a way such that the successive elements of the gradual Kesten-Tang-Yu sequence correspond covering relationships in the stable lattice under $\widetilde{P}$.
@@
Now, consider $\mu_{0}=\mathsf{IPDA}(P)$. We know from the main result of [TY14] (or [Ren22, Lemma 2]) that in 4.5, every $d$ matched to an under-demanded institution must be matched as in $\mathsf{EADAM}$. Thus, for each $h^{\prime}$ above such a $d$’s current match in $\mu_{i}$ is such that $(d,h^{\prime})$ is an illegal pair. Thus, for each step of 4.5, for each $d$ matched to an under-demanded institution, all of the institutions preferred to $d$’s current match $h$ can be removed from $d$’s list without changing the set of legal matchings. Let $P_{\mathsf{top}}$ denote the result of applying this operation successively for all steps of 4.5. Note that there are still additional illegal edges that remain on applicants’ lists in $P_{\mathsf{top}}$; since we know all matchings $\mu_{i}$ are legal (by 4.3 and the fact that $\mathsf{PN}\subseteq\mathsf{L}$), these remaining illegal pairs are some set of pairs $(d,h)$ such that $\mu_{i}(d)\prec_{d}h\prec_{d}\mu_{i+1}(d)$ for some $i$ (or such that $h\prec_{d}\mu_{1}=\mathsf{IPDA}$). However, we can see from the definition of rotations under some set of preferences (see e.g. [GI89]) that deleting any such $h$ from $d$’s list will never change whether a given rotation $\alpha$ is an actual rotation under the preferences; in the language of [GI89], a rotation which was exposed in a matching $\mu$ before any further removal of such illegal pairs must remail exposed in $\mu$ after we remove these further illegal pairs, directly by the definition of a rotation exposed in $\mu$. (Note, however, that this operation may change the predecessor relationships between the rotations, to see this, consider e.g. and type II predecessor relationships as defined in [GI89]—this predecessor relationship will be removed in the legal rotation poset.) Thus, when all illegal edges have been removed, and thus by [FZ22, Lemma 3.6] the legal set equals the stable set, we will still have that all rotations along the gradual Kesten-Tang-Yu sequence also constitute rotations in the legal lattice. Since one matching covers another in a stable matching lattice if and only if they differ by rotations, the result follows. ∎
4.1.3 Structural result
We now prove our structural result for $\mathsf{PN}$. To state this structural result, we again make heavy use of the notion of the rotation poset [GI89]. First, we introduce some notation.
Let $\mathsf{K}$ denote the sub-lattice of the legal lattice $\mathsf{L}$ which consists of all matchings that dominate $\mathsf{IPDA}$, i.e., $\mathsf{K}=\{\mu\in\mathsf{L}\ \mid\ \mu\geq\mathsf{IPDA}\}$.131313 Lattice $\mathsf{K}$ is more convenient for our purposes than $\mathsf{L}$, because in addition to maintaining the same nice structure as $\mathsf{L}$, lattice $\mathsf{K}$ has the same highest and lowest elements as $\mathsf{PN}$. This means that the lowest element of $\mathsf{PN}$, namely $\mathsf{IPDA}$ (Theorem 2.5), corresponds to the empty set of rotations in the representation of $\mathsf{K}$ in Theorem 4.7. (In contrast, the legal lattice $\mathsf{L}$ always contains the institution-side analogue of $\mathsf{EADAM}$—i.e., the matching resulting from running $\mathsf{EADAM}$ while interchanging the roles of applicants and institutions—because the definition of $\mathsf{L}$ is symmetric in the roles of applicants and institutions. Hence $\mathsf{L}$ can have a bottommost element strictly below $\mathsf{IPDA}$.) Note that $\mathsf{PN}\subseteq\mathsf{K}$ by Theorem 2.5 and Theorem 4.4. From the fact that $\mathsf{L}$ is a distributive lattice, we immediately see that $\mathsf{K}$ is a distributive lattice; hence it can be represented by a partial order on some set. We will moreover show as a simple corollary of [FZ22] a that $\mathsf{K}$ can be represented by a partial order on a set of rotations that one applies to some base matching $\mu_{0}$, exactly as for the set of stable matchings and as for $\mathsf{L}$. Our main structural result will combine 4.3 and 4.6 in order to additionally show that there is some maximal chain in $\mathsf{K}$ such that every element in this maximal chain is also in $\mathsf{PN}$.
Theorem 4.7 (Structural Result for $\mathsf{PN}$).
Fix any profile of preferences and priorities. There exists a set of rotations $R$, and a partial order on those rotations, such that $\mathsf{K}\supseteq\mathsf{PN}$ is isomorphic to the downward-closed subsets of the partial order $R$.

### v2

@@ 4.2.1 Properties of $\mathsf{PN}$
To begin, we recall one of the integral tools [Ren22] uses to prove his results for $\mathsf{PN}$ matchings. This is the notion of a (Reny-)feasible chain, which is defined as follows:
@@
A chain $\mu_{1}<\mu_{2}<\ldots<\mu_{N}$ is Reny-feasible if $\mu_{1}$ is stable, $\mu_{N}$ is Pareto-efficient, and for each $i$, $\mu_{i}$ does not violate the priority of any applicant who is Pareto-improvable in $\mu_{i-1}$. ∎
While [Ren22] reasons about Reny-feasible sequences in quite some detail, he does not prove the following fundamental fact about Reny-feasible sequences: every matching in a Reny-feasible sequence is always priority-neutral. To prove this fact, we extend the techniques from [Ren22, Lemmas 3–6].
First, we recall one simple lemma from [Ren22]:
@@
We now prove our core result regarding Reny-feasible chains:
@@
Fix some $n\in\{1,\ldots,N\}$, and suppose that $\xi$ is an improver-only violator of $\mu_{n}$. Using induction on $n^{\prime}$, we will prove the following claim: for all $n^{\prime}\in\{1,\ldots,n\}$, we have that for all applicants dd,151515[Ren22, Lemmas 4 and 5] perform related inductive arguments, but get weaker conclusions that are only sufficient to conclude that the final matching $\mu_{N}$ in the sequence is priority-neutral.
@@
To begin, we will prove using 4.6 that $\mu_{n^{\prime}}\vee\xi$ is a matching. To see this, observe that by induction, we know that for all $d\in\operatorname*{vio}(\mu_{n^{\prime}})$, we have $\mu_{n^{\prime}}(d)\succeq_{d}\xi(d)$. Thus, $\mu_{n^{\prime}}$ does not violate any $d$’s priority at $\xi(d)$. On the other hand, by the definition of an improver-only violator of $\mu_{n}$, we know that for all $d\in\operatorname*{vio}(\xi)$, we have $\xi(d)\succeq_{d}\mu_{n}(d)\succeq_{d}\mu_{n^{\prime}}(d)$. Thus, $\xi$ does not violate any $d$’s priority at $\mu_{n^{\prime}}(d)$. Thus, 4.6 implies that $\mu_{n^{\prime}}\vee\xi$ is a matching.
Now, consider any applicant $d\in\operatorname*{vio}(\mu_{n^{\prime}+1})$. By the definition of a Reny-feasible sequence, $d$ is not Pareto-improvable in $\mu_{n}^{\prime}$. Since $\mu_{n^{\prime}+1}\geq\mu_{n^{\prime}}$, also by the definition of a Reny-feasible sequence, we thus have $\mu_{n^{\prime}+1}(d)=\mu_{n^{\prime}}(d)$. Thus, since $\mu_{n^{\prime}}\vee\xi\geq\mu_{n^{\prime}}$, we must have $\mu_{n^{\prime}}\vee\xi](d)=\mu_{n^{\prime}}(d)$, and therefore $\mu_{n^{\prime}+1}(d)=\mu_{n^{\prime}}(d)\succeq_{d}\xi(d)$. This finishes the proof of the inductive step, and proves the claim for all $n^{\prime}\leq n$.
@@
We will also need the following result, which says that when looking for priority correcting adjustments, it is without loss of generality to look at matchings $\geq\mathsf{IPDA}$. We prove this result using a novel strengthening of the classical argument used to show that $\mathsf{IPDA}$ is the institution-optimal stable match, going all the way back to [GS62].
Lemma 4.8.
Consider any matching $\nu\in\mathsf{K}$, and let $\xi$ be any priority-correcting adjustment of $\nu$. Then $\xi\geq\mathsf{IPDA}$.
Moreover, consider any $\nu\geq\mathsf{IPDA}$, and suppose $\xi$ is a matching such that for all $d\in\operatorname*{vio}(\xi)$, we have $\xi(d)\succeq_{d}\nu(d)$. Then $\xi\geq\mathsf{IPDA}$.
Proof.
Recall that for each $\nu\in\mathsf{K}$, we have $\nu\geq\mathsf{IPDA}$. As in the prove of 4.7, call $\xi$ an improver-only violator of $\mu$ if $d\in\operatorname*{vio}(\xi)$ implies $\xi(d)\succeq_{d}\mu(d)$, and that priority-correcting adjustments of $\nu$ are just improver-only violators with additional requirements. Thus, the first paragraph of the lemma is actually a special case of the second paragraph.
We now prove the second paragraph of the lemma. To begin, observe that if $\xi$ satisfies the above condition for any $\nu\geq\mathsf{IPDA}$, then it also satisfies the same condition for $\nu=\mathsf{IPDA}$ (since for each $d\in\operatorname*{vio}(\xi)$, we have $\xi(d)\succeq_{d}\nu(d)\succeq_{d}\mathsf{IPDA}(d)$). Thus, it suffices to prove the lemma for the case where $\nu=\mathsf{IPDA}$. We prove the contrapositive, i.e. that if $\xi\ngeq\mathsf{IPDA}$, then there must exist a $d\in\operatorname*{vio}(\xi)$ such that $\mathsf{IPDA}(d)\succ_{d}\xi(d)$.
To this end, suppose that $\xi\ngeq\mathsf{IPDA}$. This means there exists a $d$ such that $\mathsf{IPDA}(d)\succ_{d}\xi(d)$. Let $h=\mathsf{IPDA}(d)$. If $d\succ_{h}\xi(h)$, then we are done because $d\in\operatorname*{vio}(\xi)$ and $\mathsf{IPDA}(d)\succ_{d}\xi(d)$. So suppose going forward that $\mathsf{IPDA}(h)=d\prec_{h}\xi(h)$. [^16] In what follows, we use a classical argument from [GS62] which to proves that $\mathsf{IPDA}$ is the institution-optimal stable matching, and we strengthen the conclusion of the argument. Namely, we prove that if any institution $h$ is matched to an applicant they prefer to $\mathsf{IPDA}(h)$, then not only is it the case that some applicant’s priority is violated, but the priority of some applicant who prefers $\mathsf{IPDA}$ is violated.
Now, consider some sequence of proposals and rejections made during the execution of the $\mathsf{IPDA}$ algorithm. Since $\xi(h)\succ_{h}\mathsf{IPDA}(h)$, there exists at least one $h$ such that $\xi(h)$ rejects $h$ during $\mathsf{IPDA}$. Consider the first such rejection; say, applicant $d_{1}$ rejects $h_{1}=\xi(d_{1})$ in favor of $h_{2}\neq\xi(d_{1})$; note that by the way the $\mathsf{IPDA}$ algorithm works, we have $h_{2}\succ_{d_{1}}h_{1}$. Since this is the first such rejection, we know that $h_{2}$ has not yet been rejected by $\xi(h_{2})$, so it follows that $d_{1}\succ_{h_{2}}\xi(h_{2})$. Combining these last two facts, we see that $d_{1}\in\operatorname*{vio}(\xi)$.
Moreover, by the way the $\mathsf{IPDA}$ algorithm works, $d_{1}$’s match can only improve through the remaining run of $\mathsf{IPDA}$; thus, $\mathsf{IPDA}(d_{1})\succeq_{d_{1}}h_{2}\succ_{d_{1}}h_{1}=\xi(d_{1})$. Thus, we get both that $d_{1}\in\operatorname*{vio}(\xi)$, and that $\mathsf{IPDA}(d_{1})\succ_{d_{1}}\xi(d_{1})$. This finishes the proof. ∎
@@ 4.2.2 Properties of $\mathsf{L}$
We now introduce another type of chain in order to additionally reason about $\mathsf{L}$ and $\mathsf{K}$ while bringing in the techniques of [TY14]. This chain modifies the Kesten-Tang-Yu chain (as defined in 2.3) by, informally speaking, progressing the chain with “as small of differences as possible.” To formalize this, we recall that in any lattice, an element $x$ covers $y$ if $x\geq y$, and for any $x\geq z\geq y$, we have either $x=z$ or $y=z$. Covering elements in stable lattice correspond to two matchings differing by rotations in the rotation poset. In other words, $\mu_{2}$ covers $\mu_{1}$ in a stable lattice if and only if there is a rotation $\alpha$ exposed in $\mu_{1}$ such that $\alpha\mu_{1}=\mu_{2}$.
@@
We say that a chain of matchings is a gradual Kesten-Tang-Yu chain if it can be constructed as follows: Set $P_{0}=P$ to be the full preference (and priority) profile, and let $\mu_{0}=\mathsf{IPDA}(P_{0})$. Now, for $i\geq 1$, as long as $\mu_{i-1}\neq\mathsf{APDA}(P_{i-1})$, set $P_{i}=P_{i-1}$, and set $\mu_{i}$ equal to any matching which covers $\mu_{i-1}$ in the stable matching lattice under preferences $P_{i-1}$. If $\mu_{i-1}=\mathsf{APDA}(P_{i-1})$, yet $\mu_{i-1}$ is not Pareto-optimal, then modify $P_{i}$ as in 2.3 (i.e., let $P_{i}$ denote the following operation on $P_{i-1}$: for each $d$ matched to an under-demanded institution in $\mu_{i-1}$, remove all institutions in their list strictly before $\mu_{i-1}(d)$). One can show $\mu_{i-1}$ is stable under preferences $P_{i}$ as well; we set $\mu_{i}$ to any matching stable under $P_{i}$ which covers $\mu_{i-1}$ in the stable lattice under $P_{i}$. The chain terminates with $\mu_{N}$ once $\mu_{N}$ is Pareto-optimal. ∎
By [Ren22, Lemma 2], or by the techniques of [TY14], we know that if $\mu_{i}$ in the gradual Kesten-Tang-Yu chain is not Pareto optimal, then some applicants must be matched to under-demanded institutions. [^17] In more detail, [Ren22, Lemma 2] shows that if an applicant $d$ is Pareto-improvable at $\mu_{i}=\mathsf{APDA}(P_{i})$, then $d$ cannot be matched to an under-demanded in $\mathsf{APDA}(P_{i})$. Thus, whenever $\mu_{i}=\mathsf{APDA}(P_{i})$ but $\mu_{i}$ is not Pareto-optimal, we know that $P_{i+1}\neq P_{i}$. Conversely, if the matching is Pareto-optimal, then it must equal $\mathsf{APDA}$ under the current preferences (and under the original preferences). Hence there always exists at least one gradual Kesten-Tang-Yu sequence.
Additionally, we observe that any gradual Kesten-Tang-Yu sequence is also Reny-feasible. To see why, observe that, by construction, a gradual Kesten-Tang-Yu sequence will only violate the priority of applicants matched to under-demanded institutions (and these applicants must be Pareto-unimprovable by [Ren22, Lemma 2]). Moreover, the first matching $\mu_{0}$ of a gradual Kesten-Tang-Yu chain equals $\mathsf{IPDA}$, and the last matching $\mu_{N}$ is Pareto-efficient and equals $\mathsf{EADAM}$. Thus:
Observation 4.10.
Any gradual Kesten-Tang-Yu sequence is Reny-feasible.
Hence, by 4.7, all elements of any gradual Kesten-Tang-Yu sequence are priority-neutral, and by Theorem 4.3, also legal.
@@
Fix some profile $P$ of preferences (and priorities). To begin, we recall some concepts from [FZ22]. If applicant $d$ is never matched to institution $h$ in any matching in $\mathsf{L}$, then [FZ22] call the pair $(d,h)$ an illegal pair. Let $\widetilde{P}$ denotes the altered preferences where, for all $d$, we modify $d$’s preference list by removing all $h$ such that $(d,h)$ is an illegal pair. [FZ22, Lemma 3.5] shows that the legal lattice on $P$ equals the stable lattice on $\widetilde{P}$ (and moreover that illegal pairs can always be removed from applicants preference lists without changing the set of legal matchings). Our proof proceeds by showing that $\widetilde{P}$ can be constructed from the execution of a gradual Kesten-Tang-Yu algorithm as in 4.9 (along with additional steps after the sequence) in a way such that the successive elements of the gradual Kesten-Tang-Yu sequence correspond covering relationships in the stable lattice under $\widetilde{P}$.
@@
Now, consider $\mu_{0}=\mathsf{IPDA}(P)$. We know from the main result of [TY14] (or [Ren22, Lemma 2]) that in 4.9, every $d$ matched to an under-demanded institution must be matched as in $\mathsf{EADAM}$. Thus, for each $h^{\prime}$ above such a $d$’s current match in $\mu_{i}$ is such that $(d,h^{\prime})$ is an illegal pair. Thus, for each step of 4.9, for each $d$ matched to an under-demanded institution, all of the institutions preferred to $d$’s current match $h$ can be removed from $d$’s list without changing the set of legal matchings. Let $P_{\mathsf{top}}$ denote the result of applying this operation successively for all steps of 4.9. Note that there are still additional illegal edges that remain on applicants’ lists in $P_{\mathsf{top}}$; since we know all matchings $\mu_{i}$ are legal (by 4.7 and the fact that $\mathsf{PN}\subseteq\mathsf{L}$), these remaining illegal pairs are some set of pairs $(d,h)$ such that $\mu_{i}(d)\prec_{d}h\prec_{d}\mu_{i+1}(d)$ for some $i$ (or such that $h\prec_{d}\mu_{1}=\mathsf{IPDA}$). However, we can see from the definition of rotations under some set of preferences that deleting any such $h$ from $d$’s list will never change whether a given rotation $\alpha$ is an actual rotation under the preferences: a rotation which was exposed in a matching $\mu$ before any further removal of such illegal pairs must remail exposed in $\mu$ after we remove these further illegal pairs, directly by the definition of a rotation exposed in $\mu$. (Note, however, that this operation may change the predecessor relationships; in particularly, one can show that it removes all type 2 predecessors.) Thus, when all illegal edges have been removed, and thus by [FZ22, Lemma 3.6] the legal set equals the stable set, we will still have that all rotations along the gradual Kesten-Tang-Yu sequence also constitute rotations in the legal lattice. Since one matching covers another in a stable matching lattice if and only if they differ by rotations, the result follows. ∎
4.2.3 Good Chain Property
We now prove our main structural result for $\mathsf{PN}$, which we call the “good chain property”. This result will combine 4.7 and 4.11 in order to additionally show that there is some maximal chain in $\mathsf{K}$ such that every element in this maximal chain is also in $\mathsf{PN}$.
Theorem 4.12 (Good Chain Property of $\mathsf{PN}$).